OpenStudy (anonymous):

log(11 + 1/4) without calculator.

OpenStudy (anonymous):

given log2 = 0.301 and log3= 0.477

OpenStudy (ghazi):

\[11+\frac{ 1 }{ 4 }= (11*4+1)/4=\frac{ 45 }{ 4 }\] can you do ti now?

OpenStudy (ghazi):

\[45= 5*3*3, 4=(2)^2\]

OpenStudy (anonymous):

yes but i am not given the log 5

OpenStudy (anonymous):

log(45/4)= log45-log4

hartnn (hartnn):

log 5 = log 10/2 = log 10-log 2 =1-log 2

OpenStudy (anonymous):

okay, i can do it now

OpenStudy (ghazi):

\[\log (a)^b= b \log a,... \log(a*b*c)=\log a +\log b+ \log c\]

OpenStudy (ghazi):

\[\log(5)=\log(10/2)= \log 10- \log 2=1-\log 2\]

OpenStudy (anonymous):

log(45/4) = log45-log4 = (2log3 + log5) -2log2 = 2(.477) + (1-0.310) -2(.310)

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

ok thank you! :)

OpenStudy (ghazi):

:)